OpenStudy (anonymous):

find the area y=e^x y=xe^x x=0

OpenStudy (anonymous):

Can you be more precise?

OpenStudy (anonymous):

oh. sorry. find the area enclosed by y=e^x, y=xe^x and x=0

OpenStudy (anonymous):

my guess is you are looking for \[\int_0^1e^x-xe^xdx\]

OpenStudy (anonymous):

It can also be between 0 and 10

OpenStudy (anonymous):

that is just a guess the only part that requires much work is finding where they intersect

OpenStudy (anonymous):

but, how do you know is not integral from xe^x-e^x,

OpenStudy (anonymous):

@eliassaab i don't think it can be 0 to ten because they intersect at \(x=1\)

OpenStudy (anonymous):

oh that is a good point you need to know which one is larger on the interval \([0,1]\) don't you?

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

but if \(0<x<1\) it should be fairly obvious that \(e^x>xe^x\)

OpenStudy (anonymous):

Yes, I agree. But in general you can find the area between 0 and b

OpenStudy (anonymous):

reedited question included the phrase "find the area enclosed by"

OpenStudy (anonymous):

oh, got it!!! thanks it should be from 0 to 1. also, my second question. how to get anti of xe^x?

OpenStudy (anonymous):

Integrate by parts

OpenStudy (anonymous):

i took "enclosed" to mean between the curves

OpenStudy (anonymous):

enclosed---be bounded by, i think. by part?? yo mean substitution rule?

OpenStudy (anonymous):

no not substitution, "parts"

OpenStudy (anonymous):

can you be more specific?

OpenStudy (anonymous):

\[\int udv=uv-\int vdu\]

OpenStudy (anonymous):

oh, i have not been introduced to iintegration by parts

OpenStudy (anonymous):

then i have no idea how you would do this otherwise

OpenStudy (anonymous):

i can show you how to do it in this example if you like it is a good example to start with

OpenStudy (anonymous):

oh,, yes, please, I would love to love. It is strange. We finished 6.2--volumes. my professor put this problem as final review problem. But we did not learn what you said--by parts??

OpenStudy (anonymous):

yes, "parts" is the product rule in reverse, just like "u-sub" is the chain rule in reverse

OpenStudy (anonymous):

you have \[\int xe^xdx\] and you can see that is it very easy to find the derivative of \(x\) which is 1, so it will be gone in the second step, and also easy to find the anti derivative of \(e^x\) which is just \(e^x\) so starting with \[\int xe^xdx \] and looking at \[\int udv=uv-\int vdu\]you want to make \(u=x, dv=e^xdx\) which gives \(du=dx\) and also \(v=e^x\)

OpenStudy (anonymous):

therefore \[\int xe^xdx=xe^x-\int e^xdx\]

OpenStudy (anonymous):

the now it is easy, because \[\int e^xdx=e^x\] so you have as a final result \[\int xe^xdx=xe^x-e^x\] you can check that this is right, and also why "parts" works, by taking the derivative of \[xe^x-e^x\] and seeing that indeed you get \[xe^x\]

OpenStudy (anonymous):

the product rule makes it work \[\frac{d}{dx}[xe^x-e^x]=xe^x+e^x-e^x=xe^x\] tada!

OpenStudy (anonymous):

I need a little time to think about what is "part". And I don't this is the method professor is intended. But, I think your product rule seems more straightforward.

OpenStudy (anonymous):

i am not sure what your professor intended, but i am very sure that you are not going to find the integral of \(xe^x\) without using parts

OpenStudy (anonymous):

OK, haha, Got you!!THank you so much!!

OpenStudy (anonymous):

the product rule say s \[(fg)'=fg'+gf'\] integrating both sides gives \[fg=\int fg'+\int gf'\] or \[\int fg'=fg-\int gf'\] this is usually written as \[\int udv=uv-\int vdu\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Yeah!! I completely got it !! Thanks a lot!! satellite73

OpenStudy (anonymous):

yw (again)

OpenStudy (anonymous):

haha, one more question, How do you get all the integral symbols

OpenStudy (anonymous):

latex

OpenStudy (anonymous):

if you want to see the code, right click on the equation or expression then select "show math as" then "lex commands" try it

OpenStudy (anonymous):

I just tried .And I guess I am using iPad,so it didn.t come up as you told me. But, I will try it on my laptop , now, I need to practice other problems before tomorrow's final .Thanks lot again